OpenStudy (anonymous):

When a particular cell is on open circuit, the p.d. between its terminals is 1.5 V. When a 10 Ω resistance is connected between the terminals the p.d. falls to 1.0 V and when the 10 Ω resistance is replaced by a resistance R the p.d. becomes 0.5 V. What is: a) the internal resistance of the cell; b) the value of R? thanks :)

OpenStudy (kropot72):

The current when the 10 ohm resistance is connected between the terminals is \[I=\frac{1}{10}\ amp\] The current of 1/10 amp causes a voltage drop of 0.5 V when flowing through the internal resistance of the cell. Therefore the value of internal resistance is given by \[r _{internal}=\frac{0.5}{\frac{1}{10}}=\frac{0.5}{0.1}\Omega\]